OpenStudy (anonymous):

e^-x*sqrt(3*e^2x-5/4) = e^x Solve for x?

zepdrix (zepdrix):

Hmm that's a little hard to read, lemme see if I'm looking at it correctly: \[e^{-x}\sqrt{3e^{2x-5/4}}=e^x\] is that the equation? :o

OpenStudy (anonymous):

No, the 5/4 is not part of the exponent in the square root

zepdrix (zepdrix):

\[e^{-x}\sqrt{3e^{2x}-\frac{ 5 }{ 4 }}=e^x\] Hmm ok let's see what we can figure out :)

OpenStudy (anonymous):

Exactly!!! :)

OpenStudy (anonymous):

I am stuck on how to pull the x out from the rest of the equation :'(

zepdrix (zepdrix):

Well we'll be able to use logarithms to get the x's out of the exponent, but before we do that, we have some algebra to do D:

zepdrix (zepdrix):

 Do these steps make sense so far? :o